# Nanorod Pair Complexes Manipulated via Magnetic Casimir Forces

### v1

@@ Abstract
Controlling nanoscale interactions to suppress aggregation from short-range attractive forces is a key problem in nanoengineering. Here, we demonstrate a route to modulate Casmir-Lifshitz interactions between anisotropic nanoparticles with the magnetic fluids. By semi-classical quantum electrodynamics, we study ground state dispersion forces for cylindrical dielectric nanorods made of polystyrene (PS), and zinc oxide (ZnO) embedded in toluene-based host media with gold-coated magnetite nanoparticles and also predict magnetic contributions to the non-retarded excited state interaction. The variation in magnetic permeability enables tuning between repulsive and attractive interaction and a thermally unstable and measurable magnetic Casimir traps are predicted between a pair of ZnO–PS nanoparticles whose equilibrium position can be modulated over an order of magnitude with a small variation in the size of the magnetite nanoparticle. This provides an alternative magnetic Casimir-effect pathway to reversibly tune quantum electromagnetic forces at the nanoscale for assembly and enhancement of colloidal stability.
@@
What if the very surrounding media in space itself could be engineered into a tool for nano-manipulation? For instance, can we take the background media as a magnetic fluid, in which electromagnetic fluctuations exert forces that can pull, push or most interestingly trap matters at the nanoscale level? Such fluctuation-induced interactions, which are referred to as van der Waals 1, 2 (vdW) and Casimir-Lifshitz (CL) 3, 4 forces together, typically act to bring microscopic bodies into contact, leading to unwanted self-aggregation in nanofabrication and soft-matter systems 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15. An important practical issue in nano and micro-devices is to overcome close contact attractions that causes building blocks to cluster together and an ideal solution would be to engineer systems where the interaction turns repulsive at small distances14. Beyond simply mitigating stiction, these fluctuations could also be harnessed to produce stable and controllable interaction potentials. Fundamental investigations of the influence on CL forces by a magnetic surrounding media have been largely restricted to simple geometries 16, 17, 18. Exploiting such a foundation, we have recently expanded the theoretical approach by including the contributions of transverse magnetic (TM) and transverse electric (TE) modes within a model described by the proximity force theorem 19, 20 in such a way that both energy contributions due to TM and TE-modes would partially act as candidates to enable quantum trapping between two quasi-planar surfaces 21.
@@
In this letter, we go beyond these simple geometries and study an experimentally feasible setup which comprises of pairs of cylindrical nanorods made of polystyrene (PS) and zinc oxide (ZnO). The use of the full cylindrical geometry permits a more realistic description of nanoscale behavior, albeit at the price of considerable added theoretical complexity. This stimulates further development of analytical and numerical studies of Lifshitz and vdW interaction between elongated nanoparticles 22, 23, 24, 25, 26, 27. One of us previously derived a few asymptotic limits for this kind of a system 25. Here, we examine the magnetic contributions associated with a surrounding ferrofluid for the long-range and high-temperature limits of ground-state dispersion interactions between cylindrical nanoparticles. We also predict a new kind of magnetic contribution to the retarded excited-state interaction. In an illustrative example, we study the CL interaction between a ZnO nanorod and a PS nanorod suspended in a toluene and gold-coated magnetite nanoparticle (5-7%) fluid mixtures. In the case of two unequal nanorods, we find that subtle change of nanoparticle concentration into the ferrofluid can induce a transition from short-range repulsion to long-range attraction between them. It is the fluid’s magnetic permeability and its dependence upon the magnetite nanoparticle size that dictates both interaction strength and sign, in contrast to the fluid dielectric function which remains nearly constant. Such a device could enable controlled, magnetic-induced entropic CL trapping. Early studies of CL interactions between magnetic fluida 17, 18 were mainly concentrated on the observation of a repulsive force by changing the optical properties of the medium. By contrast, we have implemented here a way in which this zero-frequency, purely entropically driven Casimir trapping is realized exclusively by tuning the fluid’s magnetic permittivity, an effect not investigated earlier.
@@
Here, we thus propose a novel experimentally feasible approach to manipulate nanorod pair complexes by suspending them in a magnetically diluted fluid, where the interplay between repulsive non-magnetic and attractive magnetic contributions to the CL force can be continuously tuned to form stable magnetic CL traps appearing as equilibria whose strength and range derive not from mechanical considerations but from spectral properties of quantum electromagnetic fluctuations.
Figure 1: Color online: Two cylindrical nanorods, one composed of ZnO with permittivity $\varepsilon_{1}$ and permeability $\mu_{1}$, and the other of PS with permittivity $\varepsilon_{2}$ and permeability $\mu_{2}$, each having an identical radius $a$, are immersed in a magnetic liquid. The magnetic liquid is composed of toluene and it contains dispersed magnetite nanoparticles of diameter $D$ and volume fraction $\phi$ and it has effective permittivity $\varepsilon_{3}$ and permeability $\mu_{3}$. The nanorods are separated by a center-to-center distance $R$.
We consider a system of two parallel cylindrical objects of radii $a$ with dielectric constant and magnetic permittivity, ($\varepsilon_{1}$, $\mu_{1}$) and ($\varepsilon_{2}$, $\mu_{2}$), respectively, embedded in a ferrofluid of dielectric constant $\varepsilon_{3}$ and permittivity $\mu_{3}$ shown in Fig. (1). The cylinders are separated by a center-to-center distance $R$. We assume that the cylinders are infinitely long in the $z$ direction, and we will work in the quasi-static limit where the wavelength of light is much larger than any other length scale in the problem. The dispersion interaction in this system is obtained using scattering approach 25, where the interaction energy per unit length (for a cylinder with length $L$) in finite temperature $T$ is (an explicit derivation is given in the Supplementary Information),
@@
where the prime in the summation means that the term for which $p=0$ must be weighted by $1/2$. The above expression is given in Matsubara frequencies $\xi_{p}=2\pi k_{B}Tp/\hbar$, where $p$ is an integer. Here, $\mathscr{D}(i\xi_{p})$ is expressed as the determinant of the scattering matrices found in the “thin cylinder approximation” e.g., the separation distance between the cylinders is much larger than their radii ($R\gg a$). We find that (see Supplementary Information for details),
@@
Here we are also interested if a dispersion interaction in excited state could be realized between two equal cylindrical nanoparticles. In order to gain an understanding into this issue, it is instructive to start by considering the corresponding analogous situation of two identical atoms; one in its excited state while the other in its ground state. In such a scenario, the interaction can be described as a superposition of a symmetric and an antisymmetric state 28. Among these, the only state that is of practical interest will be the antisymmetric one which is often very long-lived thus effectively holding the system as a trapped metastable state. As a result, a strong resonant dispersion interaction will be maintained between the atoms until (I) the atoms move sufficiently far apart or (II) a photon is emitted, which reduces most of the coherence between them. The interaction energy is due to the separation dependence of the excited energy, such that $E^{\text{res}}(R)=\hbar[\omega_{r}(R)-\omega_{r}(\infty)]$, where $\omega_{r}(R)$ represents the resonance frequency at interatomic separation $R$. The interaction energy of the excited state is 29, 30, $E^{\text{res}}(R)=s^{2}T(R|\omega_{j})\propto R^{-3},$ where $s$ is the magnitude of the transition dipole moment of the excitation and $T(R|\omega_{j})$ is the dipole–dipole coupling tensor (from the electromagnetic Green’s function) at frequency $\omega_{j}$ which accounts for how field from one oscillating dipole influences another at distance $R$ and it typically scales as $R^{-3}$ in the non-retarded limit. This can be compared with the ground state van der Waals interaction which decays with distance as $R^{-6}$31.
This analogy implies that two identical cylindrical nanoparticles can similarly form a long lasting antisymmetric excited state by virtue of a long-range resonant dispersion interaction. A corresponding expression for the retarded excited-state interaction energy, obtained semiclassically in the Supplementary Information, reads
@@
The entropic contribution ($p=0$) suggests that Eq. (S. 63) provides $R^{-3}$ dependence that is much longer ranged than the ground state interaction between two equal thin cylindrical nanoparticles that is proportional to $R^{-5}$32. It gives both magnetic and non-magnetic contributions. Notably, this previously neglected magnetic contribution in Eq. (S. 63) has the same power law as its non-magnetic contribution. However, the significance of this contribution should be judged by the low temperature/short distance non-magnetic excited state interaction. The non-retarded excited state interaction becomes attenuated by retardation due to the finite velocity of light. In the retarded limit additional effects potentially impact the interaction via real photon exchange. Ultimately, in the limit of large distances or high temperatures, the interaction is expected to be dominated by the non-magnetic and magnetic zero frequency contributions. However, it is important to emphasize that when the interaction occurs in a fluid the excited state lifetime will be significantly reduced. This will impact the potential existence of excited interactions in a fluid at large separations.
The nanoparticles we consider have large enough dimensions to approximate their effective dielectric functions with those of bulk materials. The dielectric functions at imaginary Matsubara frequencies are obtained through the Kronig–Kramer relation
@@
We use available experimental data from the literature for polystyrene (PS) 33 nanoparticles. This reference is also used to provide the dielectric function for Toluene. For zinc oxide (ZnO) 34 nanorods we use results from density functional theory (DFT) that, together with additional materials information, are described in detail in the Supplementary Information. The optical properties of the magnetic fluid are estimated by taking into account the volume fraction of commercially available gold-coated magnetite particles in a toluene mixture. We use the mixing model presented by Chen et al.35 as,
@@
Here, $\phi_{i}$ and $\varepsilon_{3,i}$ are the volume fraction and dielectric function of toluene ($i=3$), magnetite ($i=2$) and gold ($i=1$) with gold coating thickness $R_{\rm Au}$. In the current work, we consider the case where the amount of magnetite is 5% and 7% ($\phi_{2}=0.05,0.07$). Hence, it is of interest to link $\phi_{1}=\phi_{2}\times[(R_{1}/D)^{3}-1]$ where $R_{1}=D+R_{\rm Au}$. We consider four typical cases where the average diameter ($D$) of the magnetite nanoparticles is 10 nm, 13 nm, 15 nm, and 17 nm. We use experimental data from the literature for toluene 33 at $T=298.15$ K. Gold coatings are taken to have bulk properties given in the literature 36. In the main text, we report some important results that are obtained for pure magnetite particles (i.e. for $R_{1}=D$). Most of the results obtained using gold-coating are included in Sec. (S.7) of the Supplementary Information. The size of the magnetite nanoparticles and gold-coated magnetite nanoparticle can be adjusted to control trapping behaviour for pairs of unequal nanorods.
@@
Figure 2: (Color online): The parametrized dielectric functions on the imaginary axis as a function of the frequencies ${\xi}_{p}$ for PS, and magnetic toluene based on optical experiments given by van Zwol et al.33. For ZnO, we performed DFT calculation [see Sec. (S.5.1) in the Supplementary Information]. The dielectric function for the magnetic fluid is calculated using Eq. (9) and data from Ref. 36 with a volume fraction of $\phi=5\%$.
For water (close to being a non-magnetic fluid) with a fraction $\phi$ of magnetite Klimchitskaya et al. estimated the permeability to be 17, 18
@@
Toluene is only a little more magnetic ($\mu=0.99996\approx 1.00$) so within the level of accuracy in our work the permeability of both water and toluene-based mixtures can be estimated with the approximation used by Klimchitskaya and Mostepanenko for water-based mixtures 17, 18. Using Eq. (11) we derive the permeability for a range of volume fractions (from 0.05 to 0.07) and diameters (from 10 to 17 nm). The cylindrical nanoparticles we consider are non-magnetic ($\mu_{1}(0)\approx\mu_{2}(0)\approx 1$).
@@
We recently investigated the interaction between a planar PTFE surface and a spherical PS surface in the presence of a magnetic fluid 21. In this case both the non-retarded and retarded interaction could be tuned to be repulsive while zero frequency contributions to the CL interaction (both TM and TE) where attractive. This led us to conclude that it is possible to have a magnetic contribution to CL trapping 21. However, trapping still occurred when the magnetic effects were ignored (i.e. the magnetic permeability set to one).
Figure 4: (Color online): The CL force per unit length between two ZnO and PS cylindrical objects was evaluated for volume fraction $\phi=5\%$ and a nanoparticle diameter $D=17$ nm. The terms “without gold” and “with gold” refer to magnetite nanoparticles without and with a gold coating, respectively. For gold coating, layer thickness was taken to be, $R_{\rm Au}=2$ nm. Total1 is the fully retarded calculation for all frequencies without gold coating whereas Total2 reflects the retarded force with gold coating and $p=0$ denotes the zero frequency contribution.
We will in the current work explore a more versatile system consisting of nanorod pair complexes where the materials are described by the dielectric functions shown in Fig. (2). Here the intermediate magnetic fluid has a dielectric function in between those of ZnO and PS. This indicates that the non-retarded and retarded CL interaction between a ZnO and PS nanorod will be repulsive 4.
In Fig. (3) we show that in the case of a pair of ZnO and PS nanorods in a magnetic fluid with toluene mixed with 5-7% magnetite it is possible to obtain trapping that is entirely controlled by the magnetic contribution to the CL force. When the magnetic permeability of the fluid is set to one ($\mu_{3}=1$), the system only experiences repulsion. This dependence of trapping on the magnetic permeability has not been possible to achieve in the past. Notably, when the diameter of the magnetite nanoparticles decreases with fixed volume fraction of magnetite, or a weak magnetic field is applied, leading to reduced magnetic permeability of the fluid 37, the trapping occurs at increasingly larger nanorod pair separations. Ultimately, in the absence of magnetic permeability effects, there is no trapping at all and the nanorod complex is split into two separate nanorods.
@@
This framework suggests a new route for actively tuning the interacting forces between two unequal nanorods where a purely attractive potential can be driven into repulsive–attractive trapping potentials and even into regimes of complete repulsion. By a careful choice of the material parameters, one can tailor finite nanosystems that give rise to reversible adhesive behavior, stable tunable trapping and on-demand destabilization of clusters. These results demonstrate a general platform for nanocluster design in magnetic colloids where adjustable magnetic Casimir forces can be exploited for NEMS/MEMS applications.
@@
Going forward, the framework we have developed here indicates some exciting prospects for future extension of tunable dispersion-force control to biomolecular materials. In particular, these ideas could be extended to direct biopolymer complexation with artificial nanothreads via magnetic nanoparticles in biofluids with a suitable choice of materials. This has a great potential in a variety of future diagnostic and therapeutic nanomedicine research 38. It also provides a route to field-controlled assembly in hybrid soft–magnetic nanostructures and could also find utility for biosensing, targeted delivery, reconfigurable nanoarchitectures, and beyond.

### v2

@@ Abstract
Controlling nanoscale interactions to suppress aggregation from short-range attractive forces is a key problem in nanoengineering. Here, we demonstrate a route to modulate Casmir-Lifshitz interactions between anisotropic nanoparticles with the magnetic fluids. By semi-classical quantum electrodynamics, we study ground state dispersion forces for cylindrical dielectric nanorods made of polystyrene (PS), and zinc oxide (ZnO) embedded in toluene-based host media with gold-coated magnetite nanoparticles and also predict magnetic contributions to the fully retarded excited state interaction. The variation in magnetic permeability enables tuning between repulsive and attractive interaction and measurable magnetic Casimir traps are predicted between a pair of ZnO–PS nanoparticles whose equilibrium position can be modulated over an order of magnitude with a small variation in the size of the magnetite nanoparticle. This provides an alternative magnetic Casimir-effect pathway to reversibly tune quantum electromagnetic forces at the nanoscale for assembly and enhancement of colloidal stability.
@@
What if the very surrounding media in space itself could be engineered into a tool for nano-manipulation? For instance, can we take the background media as a magnetic fluid, in which electromagnetic fluctuations exert forces that can pull, push or most interestingly trap matter at the nanoscale level? Such fluctuation-induced interactions, which are referred to as van der Waals wang1928problem, Lond (vdW) and Casimir-Lifshitz (CL) casimir1948-1, Dzya forces together, typically act to bring microscopic bodies into contact, leading to unwanted self-aggregation in nanofabrication and soft-matter systems ParsegianNinham1969, dobson2006asymptotics, Munday2009, esteso2015nanolevitation, WoodsRevModPhys.88.045003, zhao2019stable, Woods2020, esteso2022effect, MundayTorque2022, MundayReview2024, Woods2024. An important practical issue in nano and micro-devices is to overcome close contact attractions that causes building blocks to cluster together and an experimental solution is to engineer systems where the interaction turns repulsive at small distances Munday2009, esteso2015nanolevitation, WoodsRevModPhys.88.045003, zhao2019stable, Woods2020, esteso2022effect, MundayReview2024, WilliamsonJPCL2025 Beyond simply mitigating stiction, these fluctuations could also be harnessed to produce stable and controllable interaction potentials. Fundamental investigations of the influence on CL forces by a magnetic surrounding media have been largely restricted to simple geometries Richmond_1971, klimchitskaya2019impact, VelichkoKlimchitskaya2020. Exploiting such a foundation, we have recently expanded the theoretical approach by including the contributions of transverse magnetic (TM) and transverse electric (TE) modes within a model described by the proximity force theorem Derjaguin1934, BLOCKI198153 in such a way that both energy contributions due to TM and TE modes would partially act as candidates to enable quantum trapping between two quasi-planar surfaces pal2025trapping.
@@
In this letter, we go beyond these simple geometries and study an experimentally feasible setup which comprises of pairs of cylindrical nanorods made of polystyrene (PS) and zinc oxide (ZnO). The use of the full cylindrical geometry permits a more realistic description of nanoscale behavior, albeit at the price of considerable added theoretical complexity. This stimulates further development of analytical and numerical studies of Lifshitz and vdW interaction between elongated nanoparticles Richmond1972, mitchell1973van, Davies1973, NylandBrevik1994, SubhojitPCCP2024, PhysRevE.111.015403. One of us previously derived a few asymptotic limits for this kind of a system NylandBrevik1994. Here, we examine the magnetic contributions associated with a surrounding ferrofluid for the long-range and high-temperature limits of ground-state dispersion interactions between cylindrical nanoparticles. We also predict a new kind of magnetic contribution to the retarded excited-state interaction. In an illustrative example, we study the CL interaction between a ZnO nanorod and a PS nanorod suspended in a toluene and gold-coated magnetite nanoparticle (5-7%) fluid mixtures. In the case of two unequal nanorods, we find that subtle change of nanoparticle concentration into the ferrofluid can induce a transition from short-range repulsion to long-range attraction between them. It is the fluid’s magnetic permeability and its dependence upon the magnetite nanoparticle size that dictates both interaction strength and sign, in contrast to the fluid dielectric function which remains nearly constant. Such a device could enable controlled, magnetic-induced entropic CL trapping. Early studies of CL interactions between magnetic fluida klimchitskaya2019impact, VelichkoKlimchitskaya2020 were mainly concentrated on the observation of a repulsive force by changing the optical properties of the medium. By contrast, we have implemented here a way in which this zero-frequency, purely entropically driven Casimir trapping is realized exclusively by tuning the fluid’s magnetic permittivity, an effect not investigated earlier.
@@
Here, we thus propose a novel experimentally feasible approach to manipulate nanorod pair complexes by suspending them in a magnetically diluted fluid, where the interplay between repulsive non-magnetic and attractive magnetic contributions to the CL force can be continuously tuned to form magnetic CL traps appearing as equilibria whose strength and range derive not from mechanical considerations but from spectral properties of quantum electromagnetic fluctuations.
Figure 1: Color online: Two cylindrical nanorods, one composed of ZnO with permittivity $\varepsilon_{1}$ and permeability $\mu_{1}$, and the other of PS with permittivity $\varepsilon_{2}$ and permeability $\mu_{2}$, each having an identical radius $a$, are immersed in a magnetic liquid. The magnetic liquid is composed of toluene and it contains dispersed gold coated (thickness $R_{\rm Au}$) magnetite nanoparticles of diameter $D$ and volume fraction $\phi$ and it has effective permittivity $\varepsilon_{3}$ and permeability $\mu_{3}$. The nanorods are separated by a center-to-center distance $R$.
We consider a system of two parallel cylindrical objects of radii $a$ with dielectric constant and magnetic permittivity, ($\varepsilon_{1}$, $\mu_{1}$) and ($\varepsilon_{2}$, $\mu_{2}$), respectively, embedded in a ferrofluid of dielectric constant $\varepsilon_{3}$ and permittivity $\mu_{3}$ shown in Fig. (1). The cylinders are separated by a center-to-center distance $R$. We assume that the cylinders are infinitely long in the $z$ direction, and we will work in the quasi-static limit where the wavelength of light is much larger than any other length scale in the problem. The dispersion interaction in this system is obtained using scattering approach NylandBrevik1994, where the interaction energy per unit length (for a cylinder with length $L$) in finite temperature $T$ is (an explicit derivation is given in Sec. (S.2) of the Supplementary Information),
@@
where the prime in the summation means that the term for which $p=0$ must be weighted by $1/2$. The above expression is given in Matsubara frequencies $\xi_{p}=2\pi k_{B}Tp/\hbar$, where $p$ is an integer. Here, $\mathscr{D}(i\xi_{p})$ is expressed as the determinant of the scattering matrix of the interacting cylinders. In general, such a matrix is infinite, but significant simplifications can be made in the “thin cylinder” approximation where the separation distance between the cylinders is much larger than their radii ($R\gg a$). In this case, terms containing Bessel functionswatson1958 with $n\geq 2$ make negligible contributions (details in Sec. (S.2) in the Supplementary Information) as they rapidly decay with $R$. Thus, we find that (see Sec. (S.2) of the Supplementary Information),
@@
Here we are also interested if a dispersion interaction in excited state could be realized between two equal cylindrical nanoparticles. The interaction energy is due to the separation dependence of the excited energy, such that $E^{\text{res}}(R)=\hbar[\omega_{r}(R)-\omega_{r}(\infty)]$, where $\omega_{r}(R)$ represents the resonance frequency at interatomic separation $R$. The interaction energy of the excited state is McLachlan, Bostrom1, $E^{\text{res}}(R)=s^{2}T(R|\omega_{j})\propto R^{-3},$ where $s$ is the magnitude of the transition dipole moment of the excitation and $T(R|\omega_{j})$ is the dipole–dipole coupling tensor (from the electromagnetic Green’s function) at frequency $\omega_{j}$ which accounts for how field from one oscillating dipole influences another at distance $R$ and it typically scales as $R^{-3}$ in the non-retarded limit. This can be compared with the ground state van der Waals interaction which decays with distance as $R^{-6}$Parsegian2006.
This analogy implies that two identical cylindrical nanoparticles can similarly form a long lasting antisymmetric excited state by virtue of a long-range resonant dispersion interaction. A corresponding expression for the retarded excited-state interaction energy, obtained semiclassically in Sec. (S.3) of the Supplementary Information, reads
@@
The entropic contribution ($p=0$) suggests that Eq. (6) provides $R^{-3}$ dependence that is much longer ranged than the ground state interaction between two equal thin cylindrical nanoparticles that is proportional to $R^{-5}$langbein1972van. It gives both magnetic and non-magnetic contributions. Notably, this previously neglected magnetic contribution in Eq. (6) has the same power law as its non-magnetic contribution. However, the significance of this contribution should be judged by the low temperature/short distance non-magnetic excited state interaction. The non-retarded excited state interaction becomes attenuated by retardation due to the finite velocity of light. In the retarded limit additional effects potentially impact the interaction via real photon exchange. Ultimately, in the limit of large distances or high temperatures, the interaction is expected to be dominated by the non-magnetic and magnetic zero frequency contributions [for more detail see Sec. (S.7) in the Supplementary Information]. However, it is important to emphasize that when the interaction occurs in a fluid the excited state lifetime will be significantly reduced. This will impact the potential existence of excited interactions in a fluid at large separations.
The nanoparticles we consider have large enough dimensions to approximate their effective dielectric functions with those of bulk materials. The dielectric functions at imaginary Matsubara frequencies are obtained through the Kramers–Kronig relation
@@
The geometries of the nanoparticles and nanorods are described within our theoretical model, whereas the material properties are macroscopic attributes. We use available room-temperature experimental data for PS as provided by van Zwol, et al.. Zwol1. For ZnO we calculate the imaginary part of the dielectric function within the density functional theory (DFT) as implemented in the Vienna Ab initio Simulation Package. VASP1999 To correctly mimic the experimental band-gap energy of 3.4 eV, we employ the hybrid functional with 37.5% Fock exact exchange, which is a widely adopted parameter choice for this oxide. The $GW$-type pseudopotentials ensure better accuracy in describing high-frequency optical transitions. Convergence with respect to $k$-space mesh is considered. Crovetto2016 Further computational details, together with additional materials information, are described in Sec. (S.5) of the Supplementary Information.
The optical properties of the magnetic fluid are estimated by taking into account the volume fraction of commercially available gold-coated magnetite particles in a toluene mixture. We use the mixing model presented by Chen et al.ChenEffPermiLayered1998 as,
@@
Here, $\phi_{i}$ and $\varepsilon_{3,i}$ are the volume fraction and dielectric function of toluene ($i=3$), magnetite ($i=2$) and gold ($i=1$) for modeling the magnetic fluid where $D$ is the diameter of the magnetite core and the thickness $R_{\rm Au}$ of the gold coating, as illustrated in Fig. (1). The radius of the gold coated nanoparticles is defined as $R_{1}=D/2+R_{\rm Au}$.
In the current work, we consider the case where the amount of magnetite is 5% and 7% ($\phi_{2}=0.05,0.07$). The corresponding volume fraction of the gold coating is therefore given by $\phi_{1}=\phi_{2}\times[(2R_{1}/D)^{3}-1]$. We consider four typical cases where the average diameter ($D$) of the magnetite nanoparticles is 10, 13, 15 and 17 nm. We use experimental data for toluene at $T=298.15$ K from the literature. Zwol1 The dielectric functions of magnetite and gold are from DFT calculations in our recent work, Carretero_PRB_2025_PhysRevB.111.085407 where room-temperature thermal electronic excitation was considered for the small-gap magnetite compound. In the main text, we report some important results that are obtained for pure magnetite particles (i.e. for $R_{1}=D/2$). Most of the results obtained using gold-coating are included in Sec. (S.8) of the Supplementary Information. The size of the magnetite nanoparticles and gold-coated magnetite nanoparticle can be adjusted to control trapping behavior for pairs of unequal nanorods.
@@
Figure 2: (Color online): The dielectric functions on the imaginary axis as functions of frequency $\xi$. The parametrized data for PS and toluene are based on optical experiments given by van Zwol et al.Zwol1. The magnetic fluid consists of a mixture of toluene and magnetite particles, using Eq. (9) for a volume fraction of $\phi=5\%$. The response functions of gold and magnetite are DFT results presented in Ref. Carretero_PRB_2025_PhysRevB.111.085407, whereas the data for ZnO are from present DFT calculations.
For water (close to being a non-magnetic fluid) with a fraction $\phi$ of magnetite Klimchitskaya et al. estimated the permeability to be klimchitskaya2019impact, VelichkoKlimchitskaya2020
@@
Toluene is only a little more magnetic ($\mu=0.99996\approx 1.00$) so within the level of accuracy in our work the permeability of both water and toluene-based mixtures can be estimated with the approximation used by Klimchitskaya and Mostepanenko for water-based mixtures klimchitskaya2019impact, VelichkoKlimchitskaya2020. Using Eq. (11) we derive the permeability for a range of volume fractions (from 0.05 to 0.07) and diameters (from 10 to 17 nm). The cylindrical nanoparticles we consider are non-magnetic ($\mu_{1}(0)\approx\mu_{2}(0)\approx 1$).
@@
We recently investigated the interaction between a planar PTFE surface and a spherical PS surface in the presence of a magnetic fluid pal2025trapping. In this case both the non-retarded and retarded interaction could be tuned to be repulsive while zero frequency contributions to the CL interaction (both TM and TE) where attractive. This led us to conclude that it is possible to have a magnetic contribution to CL trapping pal2025trapping. However, trapping still occurred when the magnetic effects were ignored (i.e. the magnetic permeability set to one).
Figure 4: (Color online): The CL force per unit length between two ZnO and PS cylindrical objects was evaluated for volume fraction $\phi=5\%,7\%$ and a nanoparticle diameter $D=17$ nm. The terms “without gold” and “with gold” refer to magnetite nanoparticles without and with a gold coating, respectively. For gold coating, layer thickness was taken to be, $R_{\rm Au}=2$ nm. Total1 is the fully retarded calculation for all frequencies without gold coating whereas Total2 reflects the retarded force with gold coating and $p=0$ denotes the zero frequency contribution.
We will in the current work explore a more versatile system consisting of nanorod pair complexes where the materials are described by the dielectric functions shown in Fig. (2). Here the intermediate magnetic fluid has a dielectric function in between those of ZnO and PS. This indicates that the non-retarded and retarded CL interaction between a ZnO and PS nanorod will be repulsive Dzya.
In Fig. (3) we show that in the case of a pair of ZnO and PS nanorods in a magnetic fluid with toluene mixed with 5-7% magnetite it is possible to obtain trapping that is entirely controlled by the magnetic contribution to the CL force. When the magnetic permeability of the fluid is set to one ($\mu_{3}=1$), the system only experiences repulsion. This dependence of trapping on the magnetic permeability has not been possible to achieve in the past. Notably, when the diameter of the magnetite nanoparticles decreases with fixed volume fraction of magnetite, or a weak magnetic field is applied, leading to reduced magnetic permeability of the fluid ZhangNature2024, the trapping occurs at increasingly larger nanorod pair separations. The trapping can also be controlled by keeping the diameter constant and varying the volume fraction. As illustrated in Fig. (5), the trend is that larger $\phi$ shifts the force minimum towards larger separations and making it more shallow. Ultimately, in the absence of magnetic permeability effects or very small volume fraction, there is no trapping at all and the nanorod complex is split into two separate nanorods.
Figure 5: (Color online:) The CL force per unit length between two cylindrical objects (ZnO and PS respectively) with the same radius $a=20$ nm as a function of center-to-center distance $R$ when immersed in a ferrofluid with different concentrations of ($\phi$=0.01-0.15) magnetite nanoparticles of diameter 17 nm.
@@
Non-magnetic Casimir forces have been exploited, both theoretically and experimentally, to produce Casimir trapping and non-adhesion, via repulsive forces, in NEMS and MEMSMunday2009, esteso2015nanolevitation, WoodsRevModPhys.88.045003, zhao2019stable, Woods2020, esteso2022effect, MundayReview2024, WilliamsonJPCL2025. We have focused in the current work on the novel results where only the zero-frequency TE term changes, as trapping occurs over a range of different $\phi$ values. The repulsion occurs when $\phi$ is tuned such that the dielectric function for the fluid is between the values of the dielectric functions of the interacting nanorods for a wide range of frequencies. Our work demonstrates that the trapping can be controlled by both changes in the size of the magnetic nanoparticles and the concentration in the fluid. The exact values for these quantities can be adjusted to achieve trapping for different separations and for different material combinations. The proposed framework suggests a novel route for actively tuning the interacting forces between two unequal nanorods where a purely attractive potential can be driven into repulsive–attractive trapping potentials and even into regimes of complete repulsion. By a careful choice of the material parameters, one can tailor finite nanosystems that give rise to reversible adhesive behavior, stable tunable trapping and on-demand destabilization of clusters. These results demonstrate a general platform for nanocluster design in magnetic colloids where adjustable magnetic Casimir forces can be exploited for NEMS/MEMS applications.
@@
Going forward, the framework we have developed here indicates some exciting prospects for future extension of tunable dispersion-force controlMunday2009, esteso2015nanolevitation, WoodsRevModPhys.88.045003, zhao2019stable, Woods2020, esteso2022effect, MundayReview2024, WilliamsonJPCL2025 to biomolecular materials. In particular, these ideas could be extended to direct biopolymer complexation with artificial nanothreads via magnetic nanoparticles in biofluids with a suitable choice of materials. This has a great potential in a variety of future diagnostic and therapeutic nanomedicine research CoeneMagneticNPJApplPhys2022. It also provides a route to field-controlled assembly in hybrid soft–magnetic nanostructures and could also find utility for biosensing, targeted delivery, reconfigurable nanoarchitectures, and beyond.